Show the equations and calculations you use for each step in the problem solution along with a brief explanation of what is being done and why it is being done for each step.
A swimming pool measures 36 ft. by 20 ft. How much fencing is needed, including material for a gate, if the fence is to be built 9 ft. from each side?

Answer by ikleyn(52778)

.
HINT  and  the  SOLUTION:

    Fencing along the longer side is 36 + 2*9 = 36 + 18 = 54 ft long.

    Fencing along the shorter side is 20 + 2*9 = 20 + 18 = 38 ft long.    

Answer.  In total,  54 + 38 + 54 + 38  ft  fencing  is  needed.